A wedge belt is required to transmit 18.5kw from an electric motor running at 1455rpm to a uniformly loaded conveyor running at 400rpm. The desired center distance is 1.4m select a suitable belt and determine the diameter of the pulleys

1
Expert's answer
2021-11-19T00:07:33-0500

P=18.5kW

"\eta_1=1455rpm"

"\eta_2=400rpm"

C=1.4m

"Speed\ ratio=\frac{1455}{400}=3.6375"

From the standard tables at speed ratio<10; service factor=1.1

Hence design power =18.5*1.1=20.35kW

From the standard tables, for P=20.35kW

The centre distance=112mm and suitable belt is flat belt